OpenStudy (anonymous):

Answer Check - What is the slope and y-intercept: 4x + 7y = 28? My answers are in the comments.

OpenStudy (anonymous):

y = 4, x = 7 y-intercept = (0, 4) Slope = 7/3

OpenStudy (godlovesme):

@Sadworld r u there?

OpenStudy (anonymous):

here

OpenStudy (godlovesme):

oh okay the y-intercept is correct but how did u get the slope?

OpenStudy (anonymous):

\[\frac{ y_2 - y_1 }{ x_2 - x_1 }; (x_1, y_1) = (7, 0); (x_2, y_2) = (0, 4)\]

OpenStudy (anonymous):

I think I got it backwards...?

OpenStudy (godlovesme):

idk lol but i think solving for y would be easier \[4x+7y=28 \] \[7y=-4x+28 \] \[y=-4/7x +4 \]

OpenStudy (anonymous):

Damn, didn't saw that one... thanks.

OpenStudy (godlovesme):

lol it happens :P np

OpenStudy (godlovesme):

thanks for the medal